Gamer Co. has no debt. Its cost of capital is 10.5 percent. Suppose the company converts to a debt–equity ratio of 1. The interest rate on the debt is 7.6 percent. Ignore taxes for this problem. What is the company’s new cost of equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of equity % What is its new WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) WACC %
Answer #1

cost of capital (wacc) that is =   10.5%
Cost of Debt =   7.60%
Debt to Equity ratio is 1. So debt = 1, Equity =1  
weight of Debt = Debt/(Debt +Equity)= 1/(1+1)=   0.5
weight of Equity= 1-0.5 =   0.5
As per MM approach formula for  
levered cost of Equity without taxes formula= (wacc - (weight of debt*Cosequitydebt))/Weight of equit  
(10.5% - (7.6%*0.5))/0.5  
13.40%  
  
So, new cost of Equity is 13.4%  
  
(b)  
As per MM approach, if there is no taxes, change in weight of debts and equity will not effect overall Required Return of assets. it means in any capital structure, wacc shall remain same.  
So new wacc shall be 10.5%  
  
Proof:   
  
WACC = (Weight of equity * Expected cost of equity)+(Weight of debt * cost on debt)  
(0.5*13.4%)+(0.5*7.6%)  
  
10.50%  

Please thumbs up